OpenStudy (anonymous):

Given y = -3x + b find the value of b if the line passes through (4,-7).

OpenStudy (anonymous):

for this problem, all you have to do is subsitute in the corrct x and y values, that are given in the point (-7)=-3(4)+b simplify this to -7=-12+b and then to 5=b

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

(: welcome